Consider two circles, a smaller one and a larger one. If the larger one has a radius that is 16 feet larger than that of the smaller circle and the ratio of the circumferences is 5:1, what are the radii of the two circles?
r1=_____________________feet
r2=_____________________feet

Answer by mananth(16949)

Radius of bigger crcle = r1
circumference C1= 2 *pi*r1
circumference of smaller circleC2 = 2*pi*r2
C/C2 = 5/1 = r1/r2
5/1 = 16/r2
r2=16/5
r2=3.2 feet